OpenStudy (anonymous):

write an equation in slope-intercept form of the line that passes through the given point and is PERPENDICULAR to the given line (4,-3); 2x-7y=14

pooja195 (pooja195):

\[\huge~\rm~2x-7y=14 \] Start by isolating y \[\huge~\rm~2x-2x-7y=-2x+14 \]

pooja195 (pooja195):

\[\huge~\rm~2x-7y=14\] Start by isolating y \[\huge~\rm~2x-2x-7y=-2x+14\] \[\frac{ \huge~\rm~-7y=-2x+14 }{ \huge~\rm-7~~~~~~~~-7 }\]